Word Problem:
A movie theater charges $7.50 admissions for and adult and $5 for a child. If 700 tickets are sold on a particular day and the total revenue received is $4500, how many tickets of each type are sold?
I need to make an equation and then solve it.

Answer by Theo(13342)

x is the number of adults.
y is the number of children.

7.5x is the total amount of revenue based on x number of adults.
5y is the total amount of revenue based on y number of children.

total number of tickets sold is 700.

equation for that is x + y = 700

total revenue is 4500.

equation for that is 7.5x + 5y = 4500

these two equations need to be solved simultensouly becsuse the same value of x and y applies to both.

equations are:

x + y = 700
7.5x + 5y = 4500

solve for y in the first equation to get y = 700 - x
replace y with 700 - x in the second equation to get 7.5x + 5y = 4500 becomes 7.5x + 5 * (700 - x) = 4500

solve for x in the second equation.

start with 7.5x + 5 * (700 - x) = 4500

simplify to get 7.5x + 5*700 - 5x = 4500

combine like terms and simplify to get 2.5x + 3500 = 4500

subtract 3500 from both sides of the equation to get 2.5x = 4500 - 3500

simplify to get 2.5x = 1000

divide both sides of the equation by 2.5 to get x = 1000 / 2.5

solve to get x = 400

x is the number of adult tickets.

since x + y = 700, then y = 300

you have 400 adult tickets and 300 children tickets.

400 + 300 = 700 so that's ok.

7.5 * 400 + 5 * 300 = 3000 + 1500 = 4500 so that's ok as well.

solution is 400 adult and 300 children tickets were sold.